# Threats, Attacks, and Defenses in Machine Unlearning: A Survey

### v3

@@ Abstract
Machine Unlearning (MU) has recently gained considerable attention due to its potential to achieve Safe AI by removing the influence of specific data from trained Machine Learning (ML) models. This process, known as knowledge removal, addresses AI governance concerns of training data such as quality, sensitivity, copyright restrictions, and obsolescence. This capability is also crucial for ensuring compliance with privacy regulations such as the Right To Be Forgotten (RTBF). Furthermore, effective knowledge removal mitigates the risk of harmful outcomes, safeguarding against biases, misinformation, and unauthorized data exploitation, thereby enhancing the safe and responsible use of AI systems. Efforts have been made to design efficient unlearning approaches, with MU services being examined for integration with existing machine learning as a service (MLaaS), allowing users to submit requests to remove specific data from the training corpus. However, recent research highlights vulnerabilities in machine unlearning systems, such as information leakage and malicious unlearning requests, that can lead to significant security and privacy concerns. Moreover, extensive research indicates that unlearning methods and prevalent attacks fulfill diverse roles within MU systems. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This underscores the intricate relationship and complex interplay among these mechanisms in maintaining system functionality and safety. This survey aims to fill the gap between the extensive number of studies on threats, attacks, and defenses in machine unlearning and the absence of a comprehensive review that categorizes their taxonomy, methods, and solutions, thus offering valuable insights for future research directions and practical implementations.
@@
Machine unlearning, threats, attacks, defensesI Introduction
@@
In AI systems, removing knowledge requires the deletion of specific data and its effects from a trained ML model. As such, machine unlearning (MU) [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] emerged as a crucial enabler of this process. In a MU system, the training dataset comprises two subsets, namely and , representing the data points to be unlearned and the remaining data points, respectively, wherein . Denote as the model trained on the entire dataset . The goal of MU is to produce an unlearned model that closely resembles the model trained solely on , i.e., .
@@
As depicted in Figure 1, an initial, yet naive, approach to implementing MU is through retraining. This method involves discarding the current model and retraining it from scratch with the remaining data after the removal of the data to be unlearned. Nevertheless, this approach can incur substantial computational expenses, making it impractical for scenarios involving large-scale models or extensive datasets. In response to these challenges, a range of efficient MU methodologies have recently been developed, focusing on either exact or approximate unlearning [19]. Exact unlearning techniques typically involve retraining but limit the scope of data involved in order to enhance efficiency compared to naive retraining approaches. On the other hand, approximate unlearning adjusts the parameters of the existing trained model to create an unlearned model that approximates one that would be obtained through retraining from scratch. This approach often requires tradeoffs between unlearning performance and efficiency. Interested readers are referred to the relevant survey papers on machine unlearning, such as [19, 10, 11, 21, 22], for more comprehensive details about MU methodologies and taxonomy. For better understanding, we first provide some definitions and previews in the subsequent part, while our contributions are also clarified.
I-A Machine Unlearning Systems
@@
Generally, as illustrated in Figure 2, the structure of machine unlearning systems designed to provide unlearning services typically includes a server and various participants. In specific, the server plays key roles in the system, including:
@@
I-B Threats, Attacks, and Defenses in MU Systems
@@
Vulnerabilities are present in machine unlearning systems, where adversaries can launch various attacks at different phases, posing significant threats to the safe deployment of unlearning services within MLaaS. Besides threats and attacks on machine unlearning systems, many studies show that unlearning and attacks serve dual roles. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This reveals a complex interaction between these elements, essential for system safety and functionality. In this section, we offer an overview of threats, attacks, and defenses in machine unlearning systems. We explore these topics in depth in Sections III, IV, and V, where we present a detailed taxonomy as shown in Figure 3.
@@
Threats. As illustrated in Figure 2, vulnerabilities are present across all phases of the unlearning system. In the training phase, for example, data contributors can maliciously craft the training data in a way that prepares for future attacks. In the unlearning phase, requesting users have the opportunity to submit malicious unlearning requests that could be designed to achieve their goals after the unlearning process. Moreover, during the post-unlearning phase, participants with access to both the trained and the unlearned models might exploit this dual access to uncover additional information leakages, thus exposing further vulnerabilities created by the unlearning process. A more detailed discussion on the various threats that may arise is available in Section III.
@@
Attacks. The threats outlined previously can be exploited to execute attacks. For instance, a data contributor might poison the training dataset, and bypass the detection by incorporating mitigation data during the training phase, which could then be removed via unlearning in the unlearning phase [25]. A requesting user could craft an unlearning request to unlearn more information than expected, thereby manipulating the system to their advantage [23]. Furthermore, an accessible user could perform enhanced membership inference attacks [26, 27, 28] by leveraging their access to both the trained and unlearned models, exploiting the differential insights provided by comparing the two models [29]. These attacks, along with threats to unlearning systems, are discussed in detail in Section III.
@@
Beyond the attacks on machine unlearning systems, extensive research has demonstrated that certain prevalent attack strategies, like backdoor attacks [30] and membership inference attacks [26], can be adapted for unlearning evaluations, such as privacy leakage audit [29], model robustness assessment [23], and proof of unlearning [31]. The analysis of attacks used for evaluating unlearning is provided in Section IV.
@@
Defenses. Potential defense methods against attacks on MU systems are explored in conjunction with attack strategies. For example, performing membership checks on data targeted for unlearning can ensure that the data specified in unlearning requests genuinely exists in the training dataset [23]. Implementing differential privacy [32, 33] can help prevent membership inference attacks, with a sacrifice of model performance [29]. Moreover, anomaly detection through model monitoring in the unlearning phase can effectively identify potential issues [34]. A corresponding discussion, along with an analysis of threats and attacks, is given in Section III.
@@
Additionally, unlearning can serve as a method of defense. It is capable of removing backdoors from trained machine learning models [35, 36, 37], and can be employed to achieve value alignment, thereby providing a defense mechanism against potential AI safety issues [38, 39, 40]. The dual role of unlearning as a defense mechanism is analyzed in Section IV.
@@
I-C Our Contributions
The increasing emphasis on security and privacy in unlearning systems has led to a clearer recognition of their vulnerabilities and the threats they face, as revealed by related studies. Moreover, the complex roles of unlearning methods and the prevalence of attacks have become more evident, underscoring their intricate interaction in maintaining the functionality and safety of MU systems. Given this backdrop, a thorough examination of existing research on threats, attacks, and defenses in MU systems is indispensable. Such a comprehensive review would not only categorize their taxonomy, methods, and solutions but also provide valuable insights for future research directions and practical implementations. Consequently, this survey is developed to address this existing gap, aiming to offer a holistic overview and critical analysis of the field.
Comparison with other surveys. Recently, several studies have focused on surveying and analyzing machine unlearning methodologies, resulting in a clear categorization of the existing survey research. These categories include general surveys on machine unlearning, highlighted by works such as [9, 19, 10, 17, 41, 11, 42, 43], which cover a wide range of topics from basic concepts and methods to the challenges faced in the field. Additionally, specific attention has been given to federated unlearning (FU) through studies like [21, 44, 45, 22]. These surveys focus on the unique issues and uses of federated unlearning compared to traditional unlearning approaches. Furthermore, potential risks and threats within federated unlearning are examined in [46]. Benchmarks and surveys on MU approaches over large language models (LLMs) are given in [47, 48, 49, 50, 39].
1.
2.
3.
We delve deeply into critical discussions concerning potential vulnerabilities and threats in unlearning systems and the challenges in defenses against these threats and attacks, highlighting key areas for future research exploration.
Organisation of the paper. The remainder of this paper is structured as follows: Section II outlines the threat models for machine unlearning systems. Building on this foundation, we conduct a thorough review of threats within the MU workflow, complete with a taxonomy, in Section III. Section IV discusses the role of unlearning in defense mechanisms, while Section V examines how attacks can serve as a means for unlearning evaluation. Section VI delves into challenges and potential directions for future research. Finally, Section VII provides a summary and concludes the paper.
II Threat Models.
II-A Attack Roles
@@ II-D Attack Phases
III Threats in Unlearning
In this section, we explore the threats inherent to machine unlearning systems. Our focus encompasses the threats of information leakage through machine unlearning, the impact of malicious unlearning practices, and other vulnerabilities that may arise within unlearning systems. We offer a thorough analysis of their methodologies based on their respective threat models (see Section II for more details) and discuss potential defense mechanisms to mitigate these risks.
III-A Information Leakage
@@ III-B Malicious Unlearning
A number of studies have successfully executed preconditioned unlearning attacks, including [34, 25, 62]. The primary distinction between [25] and [34] lies in their approaches to generating poisoned and mitigation data during the training phase. In [25], poison generation is approached as a bilevel optimization problem, as described in [63], with mitigation data produced either through label flipping or gradient matching to neutralize the poison’s effect. On the other hand, [34] outlines two methodologies for crafting the training dataset: (i) generating poisoned data by perturbing sampled clean data, with mitigation data created using the same sampled perturbed clean data but with ground-truth labels, akin to the label flipping technique in [63], and (ii) a different strategy for poisoned data generation inspired by Badnets [64], where the mitigation data contains a trigger that exerts a stronger influence on the model, hence concealing the effects of the poisoned data.
@@ III-D Summary
Post-unlearning: at this stage, defense methods here generally aim at protecting information leaked from unlearned models, or recovering the model to its state before the attack, for example, by adding differential perturbations to the unlearned models, rolling back the model, or removing backdoors for model recovery, leveraging unlearning techniques (see Section IV for more details).
@@
IV Defense through Unlearning

### v4

@@ Abstract
Machine Unlearning (MU) has recently gained considerable attention due to its potential to achieve Safe AI by removing the influence of specific data from trained Machine Learning (ML) models. This process, known as knowledge removal, addresses AI governance concerns of training data such as quality, sensitivity, copyright restrictions, and obsolescence. This capability is also crucial for ensuring compliance with privacy regulations such as the Right To Be Forgotten (RTBF). Furthermore, effective knowledge removal mitigates the risk of harmful outcomes, safeguarding against biases, misinformation, and unauthorized data exploitation, thereby enhancing the safe and responsible use of AI systems. Efforts have been made to design efficient unlearning approaches, with MU services being examined for integration with existing machine learning as a service (MLaaS), allowing users to submit requests to remove specific data from the training corpus. However, recent research highlights vulnerabilities in machine unlearning systems, such as information leakage and malicious unlearning, that can lead to significant security and privacy concerns. Moreover, extensive research indicates that unlearning methods and prevalent attacks fulfill diverse roles within MU systems. This underscores the intricate relationship and complex interplay among these mechanisms in maintaining system functionality and safety. This survey aims to fill the gap between the extensive number of studies on threats, attacks, and defenses in machine unlearning and the absence of a comprehensive review that categorizes their taxonomy, methods, and solutions, thus offering valuable insights for future research directions and practical implementations.
@@
Machine unlearning, MLaaS, threats, attacks, defensesI Introduction
@@
In AI systems, removing knowledge requires the deletion of specific data and its effects from a trained ML model. As such, machine unlearning (MU) [4] emerged as a crucial enabler of this process. As depicted in Figure 1, an initial, yet naive, approach to implementing MU is through retraining. This method involves discarding the current model and retraining it from scratch with the remaining data after the removal of the data to be unlearned. Nevertheless, this approach can incur substantial computational expenses, making it impractical for scenarios involving large-scale models or extensive datasets. In response to this challenge, a range of efficient MU methodologies have recently been developed, allowing the unlearned model to be obtained from the existing trained model without the need for retraining from scratch [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
Furthermore, with the widespread adoption of machine learning as a service (MLaaS), AI service providers are working to integrate machine unlearning functionality into their offerings to meet the growing emphasis on the RTBF in privacy regulations and the increasing need for knowledge removal from AI models. However, research indicates that vulnerabilities are present in MLaaS systems that provide unlearning services. As illustrated in Figure 2, adversarial users may launch various attacks at different stages of the MLaaS process with unlearning functionality, leading to significant security issues and privacy concerns (see detailed discussion in subsequent sections). These threats pose significant challenges to the safe deployment of machine unlearning systems.
The increasing emphasis on security and privacy in unlearning systems has led to a clearer recognition of their vulnerabilities and the threats they face, as revealed by related studies. Moreover, the complex roles of unlearning methods and the prevalence of attacks have become more evident, underscoring their intricate interaction in maintaining the functionality and safety of MU systems. Given this backdrop, a thorough examination of existing research on threats, attacks, and defenses in MU systems is indispensable. Such a comprehensive review would not only categorize their taxonomy, methods, and solutions but also provide valuable insights for future research directions and practical implementations. Thus, this survey is developed to address this existing gap, aiming to offer a holistic overview and critical analysis of the field.
Comparison with other surveys. Recently, several studies begin to study and analyze machine unlearning methodologies, resulting in a clear categorization of the existing survey research. These categories include general surveys on machine unlearning, highlighted by works such as [11, 4, 12, 19, 22, 13, 23, 24], which cover a wide range of topics from basic concepts and methods to the challenges faced in the field. Additionally, specific attention has been given to federated unlearning (FU) through studies like [25, 26, 27, 28]. These surveys focus on the unique issues and uses of federated unlearning compared to traditional unlearning approaches. Furthermore, potential risks and threats within federated unlearning are examined in [29]. Benchmarks and surveys on MU approaches over large language models (LLMs) are given in [30, 31, 32, 33, 34].
1.
2.
3.
We provide an in-depth discussion concerning potential vulnerabilities and threats in unlearning systems and the challenges in defenses against these threats and attacks, highlighting key areas for future research exploration.
Organisation of the paper. The remainder of this paper is structured as follows: Section II outlines the key concepts fundamental to this survey. Section III introduces a unified MU workflow to support the analysis and taxonomy presented in the subsequent sections. Section IV provides an overview of the threats in MU systems and the various roles of attacks and defenses. Section V provides the threat models for machine unlearning systems. Building on this foundation, we conduct a thorough review of threats within the MU workflow, complete with a taxonomy, in Section VI. Section VII discusses the role of unlearning in defense mechanisms, while Section VIII examines how attacks can serve as a means for unlearning evaluation, followed by a summary in Section IX. Section X delves into challenges and potential directions for future research. Finally, Section XI provides a summary and concludes the paper.
II Preliminaries
In this section, we provide the foundational concepts for understanding the topics covered in this survey. We introduce machine unlearning, followed by an overview of machine learning as a service. We also discuss key attacks on AI models, including poisoning attacks and membership inference attacks, to lay the groundwork for the threat and defense analysis in later sections.
II-A Machine Unlearning
In machine unlearning, as illustrated in Figure 1, the training dataset comprises two subsets, namely and , representing the data points to be unlearned and the remaining data points, respectively, wherein . Denote as the model trained on the entire dataset . The goal of MU is to produce an unlearned model that closely resembles the model trained solely on , i.e., .
Machine unlearning can be classified as either exact unlearning or approximate unlearning [4]. Exact unlearning techniques typically involve retraining but limit the scope of data involved in order to enhance efficiency compared to naive retraining approaches. On the other hand, approximate unlearning adjusts the parameters of the existing trained model to create an unlearned model that approximates one that would be obtained through retraining from scratch. This approach often requires tradeoffs between unlearning performance and efficiency. Interested readers are referred to the relevant survey papers on machine unlearning, such as [4, 12, 13, 25, 28], for more comprehensive details about MU methodologies and taxonomy.
II-B Machine Learning as a Service
Machine Learning as a Service (MLaaS) is a cloud-based service model that provides machine learning tools and infrastructure as part of a broader service offering. In an MLaaS framework, service providers host machine learning models and processing power, allowing users to submit data and receive predictions or insights without needing to develop, train, or maintain their own models [36]. As a result, MLaaS has gained widespread adoption due to its ease of use, scalability, and accessibility, enabling organizations to focus on data and application logic rather than model development and infrastructure management. Current research efforts on MLaaS include but are not limited to privacy-preserving MLaaS [37, 38] that ensure user data privacy during model training and prediction, scalability and efficiency to handle large-scale data and complex models with improved computational performance and resource management [39, 40], security and adversarial robustness to secure MLaaS systems against adversarial attacks like data poisoning and model evasion, enhancing system robustness. [41, 42].
II-C Attacks on AI Models
Attacks on AI models refer to intentional actions taken by adversaries to manipulate, exploit, or undermine the performance and integrity of machine learning systems. These attacks can occur at various stages of the AI lifecycle, including data collection, training, and deployment. Common types of attacks include adversarial attacks [43], where small, carefully crafted changes are made to input data to deceive the model into making incorrect predictions, and data poisoning attacks [44], where malicious data is introduced into the training set to degrade the model’s performance. Other threats, such as model inversion [45] and membership inference [46], aim to extract sensitive information about the training data or individuals whose data was used to train the model. Since data poisoning attacks and membership inference attacks are widely adopted in the related works reviewed in this survey, we dedicate additional space to the preliminaries of these two types of attacks to facilitate a better understanding of their mechanisms in attacks on MU systems and the defenses discussed in the subsequent sections.
Poisoning attack. Poisoning attacks are a form of attack that aims to corrupt a machine learning model by injecting malicious data into its training set [44]. These attacks are broadly categorized into untargeted and targeted poisoning attacks. Untargeted poisoning attacks aim to degrade the overall performance of the model by introducing enough corrupted data to lower its accuracy across a wide range of inputs. In contrast, in targeted poisoning attacks, the adversary focuses on manipulating the model to misclassify specific inputs or behave incorrectly in particular situations, such as causing a facial recognition system to misidentify a single person. A more specific type of targeted attack, known as a backdoor attack (BA) [47], embeds a distinct pattern or ’trigger’ into portions of the training data [47]. This trigger can be a small patch or sticker visible to humans [48], or a subtle perturbation in benign samples that is indistinguishable from human inspection [49]. Once the model is trained or fine-tuned on this compromised data, it behaves normally with standard inputs. However, when it encounters an input containing the covert trigger, the model exhibits malicious behavior aligned with the attacker’s intentions. Note that poisoning attacks and backdoor attacks can be particularly dangerous in MU systems during the unlearning process, as they can be more stealthy and difficult to detect. A detailed review of related works and further discussion on these threats will be provided in Section VI-B.
Membership inference attack. Membership inference attack (MIA) allows an adversary to obtain an attack model to determine whether a specific data point was part of the model’s training set [46]. Given access to the model’s predictions, attackers can exploit patterns or overfitting to infer if a particular individual’s data was used in training, which can lead to privacy violations. Remarkably, MIA does not require knowledge of the target model’s specific architecture or the distribution of its training data. Relying on the shadow models, a series of shadow training datasets and disjointed shadow test datasets can be synthesized to mimic the behavior of the target model so as to train the attack model. MIA can be used both to explore information leakage by analyzing the differences between the trained and unlearned models and to evaluate whether the target data has been successfully unlearned. This will be discussed in detail in Section VI and Section VIII.
III Workflow of Machine Unlearning Systems
In this section, we will define a unified workflow for MLaaS with unlearning services, which will aid in understanding the threats and attacks within these systems, as well as the corresponding defensive strategies.
@@ III Workflow of Machine Unlearning Systems
As illustrated in Figure 2, the structure of machine unlearning systems designed to provide unlearning services typically includes a server and various participants. In specific, the server plays key roles in the system, including:
@@
IV Overview of Threats, Attacks, and Defenses in MU Systems
@@
As mentioned earlier, vulnerabilities are present in machine unlearning systems, where adversaries can launch various attacks at different phases, posing significant threats to the safe deployment of unlearning services within MLaaS. Besides threats and attacks on machine unlearning systems, many studies show that unlearning and attacks serve dual roles. For instance, unlearning can act as a mechanism to recover models from backdoor attacks, while backdoor attacks themselves can serve as an evaluation metric for unlearning effectiveness. This reveals a complex interaction between these elements, which are essential for system safety and functionality. In this section, we offer an overview of threats, attacks, and defenses in machine unlearning systems. We explore these topics in depth in Sections VI, VII, and VIII, where we present a detailed taxonomy as shown in Figure 3.
@@
Threats. As illustrated in Figure 2, vulnerabilities are present across all phases of the unlearning system. In the training phase, for example, data contributors can maliciously craft the training data in a way that prepares for future attacks. In the unlearning phase, requesting users have the opportunity to submit malicious unlearning requests that could be designed to achieve their goals after the unlearning process. Moreover, during the post-unlearning phase, participants with access to both the trained and the unlearned models might exploit this dual access to uncover additional information leakages, thus exposing further vulnerabilities created by the unlearning process. A more detailed discussion on the various threats that may arise is available in Section VI.
@@
Attacks. The threats outlined previously can be exploited to execute attacks. For instance, a data contributor might poison the training dataset, and bypass the detection by incorporating mitigation data during the training phase, which could then be removed via unlearning in the unlearning phase [52]. A requesting user could craft an unlearning request to unlearn more information than expected, thereby manipulating the system to their advantage [50]. Furthermore, an accessible user could perform enhanced membership inference attacks [46, 53, 54] by leveraging their access to both the trained and unlearned models, exploiting the differential insights provided by comparing the two models [55]. These attacks, along with threats to unlearning systems, are discussed in detail in Section VI.
@@
Beyond the attacks on machine unlearning systems, extensive research has demonstrated that certain prevalent attack strategies, like backdoor attacks [56] and membership inference attacks [46], can be adapted for unlearning evaluations, such as privacy leakage audit [55], model robustness assessment [50], and proof of unlearning [57]. The analysis of attacks used for evaluating unlearning is provided in Section VII.
@@
Defenses. Potential defense methods against attacks on MU systems are explored in conjunction with attack strategies. For example, performing membership checks on data targeted for unlearning can ensure that the data specified in unlearning requests genuinely exists in the training dataset [50]. Implementing differential privacy [58, 59] can help prevent membership inference attacks, with a sacrifice of model performance [55]. Moreover, anomaly detection through model monitoring in the unlearning phase can effectively identify potential issues [60]. A corresponding discussion, along with an analysis of threats and attacks, is given in Section VI.
@@
Additionally, unlearning can serve as a method of defense. It is capable of removing backdoors from trained machine learning models [61, 62, 63], and can be employed to achieve value alignment, thereby providing a defense mechanism against potential AI safety issues [64, 34, 65]. The dual role of unlearning as a defense mechanism is analyzed in Section VII.
@@
V Threat Models
In this section, we define the threat models based on our proposed unified MU workflow described in Figure 2, regarding the role of attackers in unlearning services, the attackers’ goals, their knowledge of the MU system, and the phases during which attacks occur. This analysis of threat models will provide a deeper understanding of the existing related works discussed in the subsequent sections.
V-A Attack Roles
@@ V-D Attack Phases
VI Threats in Unlearning
In this section, we explore the threats inherent to machine unlearning systems. Our focus encompasses the threats of information leakage through machine unlearning, the impact of malicious unlearning practices, and other vulnerabilities that may arise within unlearning systems. We offer a thorough analysis of their methodologies based on their respective threat models (see Section V for more details) and discuss potential defense mechanisms to mitigate these risks.
VI-A Information Leakage
@@ VI-B Malicious Unlearning
A number of studies have successfully executed preconditioned unlearning attacks, including [60, 52, 77]. The primary distinction between [52] and [60] lies in their approaches to generating poisoned and mitigation data during the training phase. In [52], poison generation is approached as a bilevel optimization problem, as described in [78], with mitigation data produced either through label flipping or gradient matching to neutralize the poison’s effect. On the other hand, [60] outlines two methodologies for crafting the training dataset: (i) generating poisoned data by perturbing sampled clean data, with mitigation data created using the same sampled perturbed clean data but with ground-truth labels, akin to the label flipping technique in [78], and (ii) a different strategy for poisoned data generation inspired by Badnets [79], where the mitigation data contains a trigger that exerts a stronger influence on the model, hence concealing the effects of the poisoned data. Camouflaged sample generation based on data influence is explored in [80], along with a comprehensive performance evaluation of various machine unlearning methods.
@@ VI-D Summary
Post-unlearning: at this stage, defense methods here generally aim at protecting information leaked from unlearned models, or recovering the model to its state before the attack, for example, by adding differential perturbations to the unlearned models, rolling back the model, or removing backdoors for model recovery, leveraging unlearning techniques (see Section VII for more details).
@@
VII Defense through Unlearning